### v5

# LOLA – An Open-Source Massively Multilingual Large Language Model

Abstract
This paper presents LOLA, a massively multilingual large language model trained on more than 160 languages using a sparse Mixture-of-Experts Transformer architecture. Our architectural and implementation choices address the challenge of harnessing linguistic diversity while maintaining efficiency and avoiding the common pitfalls of multilinguality. Our analysis of the evaluation results shows competitive performance in natural language generation and understanding tasks. Additionally, we demonstrate how the learned expert-routing mechanism exploits implicit phylogenetic linguistic patterns to potentially alleviate the curse of multilinguality. We provide an in-depth look at the training process, an analysis of the datasets, and a balanced exploration of the model’s strengths and limitations. As an open-source model, LOLA promotes reproducibility and serves as a robust foundation for future research. Our findings enable the development of compute-efficient multilingual models with strong, scalable performance across languages.
LOLA – An Open-Source Massively Multilingual Large Language Model
Nikit Srivastava1, Denis Kuchelev, Tatiana Moteu Ngoli1, Kshitij Shetty2, Michael Röder1, Hamada M. Zahera1, Diego Moussallem1, Axel-Cyrille Ngonga Ngomo1 Data Science Group, Paderborn University, Germany 1 {nikit.srivastava, tatiana.moteu, michael.roeder, hamada.zahera, diego.moussallem, axel.ngonga}@upb.de 2 kshitij@mail.upb.de
1 Introduction
Large Language Models (LLMs) have shown tremendous capability across a diverse set of tasks in recent years Radford et al. (2019); Kaddour et al. (2023). This progress has propelled research, with many chat-based LLMs 111 ChatGPT: chat.openai.com; LLAMA: llama.meta.com; Mistral: mistral.ai; Gemini: gemini.google.com; Deepseek: deepseek.com. gaining popularity among general users. However, concerns remain, particularly regarding their accessibility for multilingual usage Joshi et al. (2020) and open-source licensing policies Liu et al. (2023). The number of competent LLMs significantly decreases for languages other than English Üstün et al. (2024). This, combined with the curse of multilinguality—a phenomenon in which the ability of models to generalize across multiple languages diminishes unless their capacity is significantly expanded Conneau et al. (2020)—means non-English speakers often have access to inferior systems. Additionally, many new models Jiang et al. (2023); Achiam et al. (2024); Dubey et al. (2024) are pay-to-use, require personal information, or do not fully disclose training details, creating significant hurdles for multilingual research.
To advance multilingual language modeling, we introduce LOLA,222 Source Code: github.com/dice-group/LOLA; Model Weights: huggingface.co/dice-research/lola_v1. a massively multilingual model that follows a GPT-style Radford et al. (2019) decoder-only architecture with sparse Mixture-of-Experts (MoE)layers Shazeer et al. (2017). MoE architectures have shown strong performance on learning the underlying structure of the data  Chen et al. (2022) but their application in multilingual LLMs remains underexplored. MoE models can effectively increase model capacity with minimal additional computational cost, offering the possibility of leveraging implicit clusters like language family groups and playing a crucial role in addressing the challenges of multilinguality.
Language family groups, consisting of languages sharing common ancestral roots, offer opportunities for enhancing language models. Despite linguistic diversity, these families exhibit structural, syntactic, and semantic similarities Rowe and Levine (2015) that can be exploited to improve performance across related languages. Our goal is to leverage MoE’s strengths to exploit the phylogenetic structure of languages and achieve better prediction performance. In particular, the shared and non-shared parameters of MoE-based models offer a promising approach to mitigating the curse of multilinguality by increasing capacity while remaining compute efficient Shazeer et al. (2017). By exploiting language families, we aim to close gaps in current models—particularly for low-resource languages—by enhancing cross-linguistic transfer learning.
Another important factor is the availability of LLMs as a free resource, accessible for anyone to use, modify, and redistribute without discrimination against any individuals or purposes. Many popular LLMs that claim to be "open source" either withhold their training datasets (e.g., Mistral, Grok333github.com/xai-org/grok-1), fail to publish their training code (e.g., Llama, Grok), or do not release their inference code (e.g., Grok-2444x.ai/blog/grok-2Spectrum (2024). In some cases, these models are released under licenses that are restrictive, discriminatory, or impose additional conditions Liesenfeld et al. (2023); Liesenfeld and Dingemanse (2024). The artifacts and components used in LOLA were selected based on their suitability for training massively multilingual LLMs while minimizing licensing concerns. All chosen components are obtainable, modifiable, and redistributable in accordance with the terms of their original licenses.
To assess LOLA’s performance, we evaluated it on four task types: 1. Question Answering (Q&A), 2. Reasoning, 3. Natural Language Inference (NLI), and 4. Reading Comprehension. In total, we assessed the model across 13 multilingual tasks, comparing it to 17 other models grouped into three categories based on their active parameter count. [^5] The number of parameters a model utilizes per token Fedus et al. (2022). This distinction is crucial for understanding the efficiency and performance of MoE models. Our results demonstrate strong performance across most tasks, though we note the limitations in 1. tasks involving factual and mathematical Q&A; and 2. comparisons with models that use more than five times the active parameters of LOLA. These findings are discussed in detail later in the paper.
Beyond presenting the multilingual model as our main contribution, we address the following key research questions:
  1. 1.
    Does training a Mixture-of-Experts model on a wide variety of languages enhance generalization or lead to confusion?
  2. 2.
    How do experts impact the model’s capacity to leverage implicit language groups?
  3. 3.
    What are the potential limitations?
2 Related Work
The development of LLMs has gained significant momentum since the introduction of the Transformer architecture by Vaswani et al. (2017). As LLMs grew in size and complexity, their capacity to model increasingly nuanced linguistic patterns expanded. Models like GPT3 and Llama Brown et al. (2020); Touvron et al. (2023) showcased the ability of large models to perform few-shot learning, a significant milestone that further highlighted the flexibility of Transformer-based architectures. As the need to extend their capabilities to handle multiple languages effectively became increasingly apparent, research into multilingual LLMs surged, aiming to enable performance across diverse languages with a single model, reducing the need for language-specific systems Zhu et al. (2024). Key efforts in this area include systems such as mBERT, XLM-R, mT5, and BLOOM Devlin et al. (2019); Conneau et al. (2020); Xue et al. (2021); Scao et al. (2022), with more recent models like Tower, SeaLLM, and Breeze Alves et al. (2024); Nguyen et al. (2024b); Hsu et al. (2024) focusing on adapting primarily English-pretrained models into multilingual ones through continued training. However, research in multilingual LLMs faces several challenges, particularly in balancing performance across languages while keeping training costs manageable, as emphasized by Conneau et al. (2020).
One of the significant challenges with scaling LLMs is the computational cost associated with training and deploying models with billions or trillions of parameters. To address this, the Mixture-of-Experts (MoE)paradigm has emerged as a promising approach for efficiently scaling large models. The MoE architecture proposed by Shazeer et al. (2017) introduces the concept of sparsity, where only a subset of the model’s parameters is activated during each forward pass, thereby reducing the computational burden while maintaining high performance. Their approach demonstrated that models could achieve state-of-the-art performance while being computationally efficient. Later approaches, such as GShard and Switch Transformers Lepikhin et al. (2021); Fedus et al. (2022), extended the MoE framework by simplifying routing and enhancing scalability, enabling models with over a trillion parameters while maintaining efficient computational costs and setting new benchmarks in large-scale model training. These advances led to increased research in MoE-based LLMs, resulting in models like GLaM, DeepSpeed MoE and Mixtral Du et al. (2022); Rajbhandari et al. (2022); Jiang et al. (2024).
Given the unique architecture of the MoE-based LLMs, Machine Translation (MT)models have explored its potential in language grouping. Several MT systems, such as M2M, NLLB, and Lingual-SMoE Fan et al. (2021); Team et al. (2022); Zhao et al. (2024), have trained MoE-based models to enable many-to-many translation, leveraging either learned or custom expert-routing mechanisms that assigns experts based on the language. Systems like NLLB continue to demonstrate state-of-the-art MT performance to this day Zhu et al. (2024). In the case of pre-trained base models, Zoph et al. (2022) briefly touch upon the multilingual nature of MoE models, though they primarily note that expert load balancing loss constrains the model’s capacity to assign language-specific experts. Despite these advances, the application of MoE for pre-training massively multilingual LLMs remains underexplored. This research contributes to addressing that gap.
3 Model Overview
Figure 1: Three-level overview of the LOLA architecture. The left-most block provides a high-level overview of the layers within LOLA, including the alternating standard and Mixture-of-Experts (MoE)-based decoder blocks. The middle block gives a detailed view of the MoE-based decoder block structure. The right-most block zooms in on the inner workings of each MoE layer, showing how the top-1 gating mechanism selects from multiple expert FFNs.
Our model is based on a GPT-style Radford et al. (2019) decoder-only Transformer architecture Vaswani et al. (2017). We replace the standard feed-forward layers (FFNs) with Mixture-of-Experts (MoE) layers in every alternate Transformer layer. These MoE layers utilize a top-1 gating mechanism inspired by the Switch Transformer Fedus et al. (2022) due to its simplicity and effectiveness. The architecture consists of 24 decoder layers with a model hidden and embedding dimension of 2048, 16 attention heads, a maximum sequence length of 2048, and each MoE layer includes 16 experts. We use the GELU Hendrycks and Gimpel (2017) non-linearities and the Adam Kingma and Ba (2015) optimizer for our model. Based on this configuration, our model has 1.3 billion active parameters out of 7.4 billion total parameters. Due to this sparsity, our model has a training/inference cost similar to that of a 1.3 billion dense model. [^6] Number of parameters activated in a single forward and backward pass. Figure 1 provides a multi-level overview of the model architecture. The model configuration and training are facilitated using the Megatron-DeepSpeed777github.com/microsoft/Megatron-DeepSpeed framework, which is based on Shoeybi et al. (2020); Rajbhandari et al. (2022).
3.1 Routing Mechanism in MoE Layers
For routing tokens through the MoE layers with $N$ (i.e., 16) experts, we first compute the logits for the gating function. These logits are then passed through a Softmax function to calculate the probability for each expert:
$h(x)=W_{g}\cdot x,$ (1)
$G_{i}(x)=\frac{\exp(h(x)_{i})}{\sum_{j=1}^{N}\exp(h(x)_{j})},$ (2)
where $h(x)$ contains the logit vectors for all experts, $W_{g}$ is the gating weight matrix, and $x$ is the input. The logit vector and gating probability of the $i$-th expert is denoted by $h(x)_{i}$ and $G_{i}(x)$ respectively.
Once the gating probabilities are computed, the output of the MoE layer is calculated by selecting the most probable expert $i^{*}$ and multiplying its gating probability $G_{i^{*}}(x)$ with the output of the corresponding expert $E_{i^{*}}(x)$:
$i^{*}=\operatorname*{arg\,max}_{i}G_{i}(x),$ (3)
$\text{MoE}(x)=G_{i^{*}}(x)\cdot E_{i^{*}}(x).$ (4)
3.2 Training and Loss Functions
Our model is pre-trained using a causal language modeling task Radford and Narasimhan (2018), where the objective is to minimize the cross-entropy loss alongside an auxiliary MoE loss.This auxiliary loss, inspired by works such as Shazeer et al. (2017), Lepikhin et al. (2021), and Fedus et al. (2022), is used to ensure stable training and effective load balancing among the experts. The auxiliary loss incorporates two vectors:
  • $P$ represents the average weight assigned to all tokens for each expert.
  • $f$ denotes the fraction of tokens allocated to each expert.
Given an input sequence $S=\{s_{1},s_{2},s_{3},\dots,s_{T}\}$ of length $T$, and $N$ experts in each MoE layer, for each expert $i=1,2,\dots,N$, these vectors are defined as:
$P_{i}=\frac{1}{T}\cdot\sum_{t=1}^{T}G_{i}(s_{t}),$ (5)
$f_{i}=\frac{1}{T}\cdot\sum_{t=1}^{T}M_{i}(s_{t}),$ (6)
where:
  • $G_{i}(s_{t})$ is the gating weight assigned to expert $i$ for token $s_{t}$,
  • $M_{i}(s_{t})$ is a binary mask indicating whether token $s_{t}$ is routed to expert $i$, determined by the top-1 gating mechanism Shazeer et al. (2017).
The auxiliary loss $l_{\text{aux}}$ is formulated as:
$l_{\text{aux}}=N\cdot\sum_{i=1}^{N}P_{i}\cdot f_{i},$ (7)
which represents the scaled dot product between $P$ and $f$.
For the language modeling task, the cross-entropy loss is computed as:
$\mathcal{L}_{\text{CE}}=-\frac{1}{T}\cdot\sum_{t=1}^{T}\log p(s_{t}\mid s_{<t}).$ (8)
The final loss function for the model combines the cross-entropy loss and the auxiliary loss:
$\mathcal{L}_{\text{final}}=\mathcal{L}_{\text{CE}}+\alpha\cdot l_{\text{aux}},$ (9)
where $\alpha$ is the multiplicative coefficient for the auxiliary loss. Throughout this work, we set $\alpha=10^{-2}$ based on the recommendations by Fedus et al. (2022).
3.3 Training Data and Setup
The model was trained on data sampled from the CulturaX Nguyen et al. (2024a) dataset, which consists of raw text documents in 167 languages, amounting to over 6 trillion tokens from more than 7 billion documents (see Appendix A.6 for train sample details). We tokenized the data using the SentencePiece Kudo and Richardson (2018) tokenizer with a vocabulary size of 100,000.
Training was conducted on 96 NVIDIA A100 GPUs [^8] GPU Model: NVIDIA A100-SXM4-40GB with a total compute of approximately 44,000 GPU hours. The model was trained for 19 days, consuming a total of 465 billion tokens across a batch size of 768 documents. [^9] Further training details in Appendix A.2
4 Evaluation
Model Params (B) Consumed Tokens (T) Max Seq. Length Languages Category
Glot500m Imani et al. (2023) 0.39 - 512 500 1
XLM-R Large Conneau et al. (2020) 0.55 6 512 100 1
mBART Liu et al. (2020) 0.68 1.8 1024 25 1
BLOOM-1B1 Scao et al. (2022) 1.10 0.341 Arbitrary 48 1
MT5 Large Xue et al. (2021) 1.20 1 Arbitrary 101 1
mGPT Shliazhko et al. (2024) 1.30 0.440 2048 61 1
BLOOM-1B7 Scao et al. (2022) 1.70 0.341 Arbitrary 48 1
XLM-R XL Conneau et al. (2020) 3.50 6 Arbitrary 100 2
MT5 XL Xue et al. (2021) 3.70 1 Arbitrary 101 2
UMT5 XL Chung et al. (2023) 3.70 1 Arbitrary 107 2
TowerBase 7B Alves et al. (2024) 6.74 2 Arbitrary 10 3
Mistral v0.3 Jiang et al. (2023) 7.00 - 32768 5 3
Falcon Almazrouei et al. (2023) 7.00 1.5 2048 2 3
BLOOM-7B1 Scao et al. (2022) 7.10 0.366 Arbitrary 48 3
SeaLLM v2 Nguyen et al. (2024b) 7.38 - Arbitrary 10 3
SeaLLM v2.5 Nguyen et al. (2024b) 7.38 - Arbitrary 10 3
Breeze Hsu et al. (2024) 7.49 - Arbitrary 2 3
LOLA (Our Model) 1.3 0.465 2048 167 1
Table 1: Characteristics of models used for comparison in the evaluation, including model names, active parameter sizes (in billions), the number of consumed tokens (in trillions), maximum sequence length, and the number of languages each model was trained on. The models are grouped by their size categories (see appendix Figure 4).
4.1 Models
After reviewing the available multilingual LLMs, we selected 17 models with active parameters ranging from 300 million to 7.5 billion. Table 1 provides a list of the selected models along with further details. The selection was based on the following criteria: 1. They are base pretrained models without any fine-tuning; 2. The weights are openly accessible without requiring personal information beyond name and email; 3. Model weights are available via Huggingface101010huggingface.co  [^11] Required for the evaluation framework.; 4. The models are compatible with our evaluation hardware setup. [^12] Single NVIDIA A100 with 40GB GPU memory, 100GB of CPU memory, and 16 CPU cores. Given the wide range of active parameters, we decided to group the models based on their sizes. We employ the distortion [^13] Mean sum of squared distances to centers. and silhouette [^14] Mean ratio of intra-cluster and nearest-cluster. scores to determine the optimal number of categories, which was identified as 3 (see Appendix A.1). Subsequently, K-Means clustering was used to classify the models into 3 categories (1-3). Although LOLA falls within Category-1, we compare and analyze its performance against each category.
4.2 Tasks
We evaluate LOLA on $13$ multilingual benchmarks datasets/tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), MMLU Hendrycks et al. (2021), MGSM Direct and MGSM Native CoT Shi et al. (2022), PAWS-X Yang et al. (2019), TruthfulQA Lin et al. (2022a), XCOPA Ponti et al. (2020), XNLI Conneau et al. (2018), XStoryCloze Lin et al. (2022b), XWinograd Tikhonov and Ryabinin (2021), and Belebele Bandarkar et al. (2023). We use the multilingual versions of originally English tasks (ARC, HellaSwag, MMLU, and TruthfulQA) introduced in OKAPI by Dac Lai et al. (2023). Details of these evaluation tasks are provided in Table 2. We utilize the Language Model Evaluation Harness framework by Gao et al. (2024) for evaluations. Examples from these tasks can be found in Appendix A.3.
Type Task Languages
Q&A ARC 31
MGSM (Direct) 11
MGSM (Native CoT) 11
TruthfulQA 31
MMLU 34
Reasoning HellaSwag 30
XCOPA 11
XStoryCloze 11
XWinograd 6
NLI PAWS-X 7
XNLI 15
Reading LAMBADA 5
Comprehension Belebele 122
Table 2: Evaluation tasks used to evaluate LOLA, along with the number of languages covered by each task.
We group the tasks into four main categories: 1. Question Answering (Q&A) 2. Reasoning, 3. Natural Language Inference (NLI), and 4. Reading Comprehension. We briefly describe each category and the corresponding tasks below:
4.2.1 Question Answering (Q&A)
This category includes tasks that require knowledge across various domains such as mathematics, philosophy, law, and medicine. ARC is a multiple-choice science question dataset for grades 3 to 9, requiring reasoning Clark et al. (2018). MGSM is a benchmark of grade-school math problems requiring multi-step reasoning, with two variations: MGSM (Direct) and MGSM (Native CoT), the latter including Chain-of-Thought prompts in the target language [^15] The target language for model evaluation. Shi et al. (2022). TruthfulQA measures a model’s ability to generate truthful answers to factual questions Lin et al. (2022a). MMLU is a large-scale multitask benchmark of multiple-choice questions spanning a wide range of topics Hendrycks et al. (2021).
4.2.2 Reasoning
This category includes tasks that require commonsense reasoning. HellaSwag assesses a model’s commonsense reasoning capabilities Zellers et al. (2019). XCOPA evaluates a model’s ability to transfer commonsense reasoning across multiple languages Ponti et al. (2020). XStoryCloze tests understanding of everyday situations through causal and relational information in daily events Lin et al. (2022b). XWinograd is a multilingual version of the Winograd Schema Challenge, requiring resolution of ambiguities in sentences differing by only one or two words, necessitating world knowledge and complex reasoning Tikhonov and Ryabinin (2021).
4.2.3 Natural Language Inference (NLI)
This category assesses the ability to identify relationships between sentences, such as paraphrasing and textual entailment. PAWS-X contains challenging paraphrase identification pairs derived from Wikipedia and Quora Yang et al. (2019). XNLI evaluates cross-lingual sentence representations by testing textual entailment Conneau et al. (2018).
4.2.4 Reading Comprehension
This category assesses reading comprehension abilities, requiring models to predict the next word or select the correct answer from given options. LAMBADA evaluates a model’s text understanding through word prediction Paperno et al. (2016). Belebele is a multilingual reading comprehension dataset evaluating models on languages with varying resource levels (high, medium, and low) Bandarkar et al. (2023).
4.3 Performance Metrics
As evaluation metrics, we employ the following:
Accuracy is a metric that assesses how frequently an input is predicted by the model to be the correct class. It is calculated by computing the ratio of correctly predicted instances to the total number of instances. This metric is used by all evaluation tasks except MGSM.
Exact Match measures the match between a reference and predicted parameter. It sums the exact match scores (1 for an exact match, 0 otherwise) and divides by the total number of predictions. This metric is used only for MGSM tasks, utilizing the flexible-extract implementation by Gao et al. (2024) to account for formatting differences.
4.4 Results
We configure our experiments based on each distinct combination of task, model, language, and the number of shots for few-shot learning. The shot settings include zero-shot, one-shot, and few-shot (i.e., 5). Altogether, we perform over 14,000 unique experiments. Given the extensive scale of these experiments, the results are not included directly in the main text for brevity. Instead, information and links to the detailed result tables are provided in Appendix A.4. A comprehensive analysis and discussion over these results is presented in the subsequent section.
5 Analysis
We present our analysis of LOLA in two subsections. In the first subsection, we discuss our key insights derived from the evaluation results. Next, we analyze LOLA’s learned MoE routing, focusing on its ability to leverage language family groupings, which aligns with our core motivation and intuition behind MoE for multilingual LLMs.
5.1 Result Analysis
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 2: Comparison of LOLA’s zero-, one- and few-shot performance against the other multilingual models across all supported combinations of tasks and languages, categorized by model size. The left side shows the results from the Wilcoxon signed-rank test, indicating whether LOLA significantly outperforms (Wins), shows no significant difference (Inconclusive) or is outperformed by (Losses) other models. On the right is the average performance comparison to confirm whether LOLA is on average better than (Wins), the same as (Ties), or worse than (Losses) the other models.
We assess LOLA’s performance relative to other models by evaluating the results across all languages for each task, employing two methods: 1. using the Wilcoxon signed-rank test Wilcoxon (1945) to determine the statistical significance of differences between performance distributions (with a $p$-value threshold of $0.05$); and 2. comparing average performance across all languages to provide a simplified overview.
These comparisons allow us to examine LOLA’s performance across various levels of granularity, including: 1. the model’s overall performance against all other models on the full set of tasks and languages; 2. its performance on specific task types; and 3. its performance on individual tasks. For brevity, we discuss the model’s overall performance in this subsection, with more detailed analyses provided in Appendix A.5.
Figure 2 shows that LOLA consistently outperforms Category-1 and Category-2 models but underperforms relative to Category-3 models, which are at least five times larger (see Table 1). Nonetheless, LOLA’s strong performance against Category-2 models—on average 2.8 times larger and trained on twice as many tokens—highlights its efficiency in multilingual settings with a substantially smaller computational footprint.
To summarize the finer granularity levels (Appendix A.5), we derive the following additional key insights about LOLA’s performance:
Strengths: 1. strong performance in NLI, Reasoning, and Reading Comprehension tasks; and 2. competitiveness with Category-3 models in NLI tasks.
Weaknesses: 1. limited gains on Q&A tasks, with particularly poor performance on MGSM; and 2. inferior few-shot performance compared to zero- and one-shot settings.
While the model’s strengths can be attributed to its generalization capabilities, its weaknesses may be due to several factors. The subpar Q&A performance may stem from LOLA’s limited factual grounding due to restricted training data per language Fierro and Søgaard (2022). Furthermore, the challenges on MGSM are likely due to the lack of a specialized tokenizer for arithmetic data and the absence of coding and LATEX data during training Yuan et al. (2023). The diminished few-shot performance may be caused by the model’s 2048-token sequence limit, which truncates essential context.161616 During evaluation overflowing sequences are truncated from the left.
These findings contribute to answering our first research question: Does training a Mixture-of-Experts model on a wide variety of languages enhance generalization or lead to confusion? The results indicate that training across diverse languages enhances generalization, particularly in NLI, Reasoning, and Reading Comprehension tasks; challenges persist in Q&A tasks, which may necessitate additional data or specialized pre-training.
5.2 MoE Analysis
In this subsection, we discuss our second research question: How do experts impact the model’s capacity to leverage implicit language groups?
We answer this question by analyzing whether there is a correlation between the activity of the experts within the model and groups of languages that share common features. To this end, we measure the activation of the experts on all layers across 106 languages. [^17] Languages for which CulturaX has at least 10,000 documents. Based on these activities, we create a vector for each language comprising the activation of the experts when processing documents of this language. Based on these vectors, we calculate a language-to-language distance matrix using the normalized Euclidean distance. We compare our distance matrix with distance matrices of the URIEL project Littell et al. (2017) comprising pairwise language distances based on a variety of features like 1. their syntactic features, 2. their phonological features, 3. their geographical location, and 4. their position in the Glottolog tree of language families Hammarström et al. (2015). We calculate the Pearson correlation coefficients between these matrices and our matrix. Our results indicate a weak positive linear correlation between the activity of our model’s experts and the distance of the languages within the language family tree. This correlation grows stronger when we focus the analysis on those languages for which the model saw more training documents, up to a correlation of 0.55 for the 23 languages that have at least 1 million documents in our training data. [^18] The Pearson correlation values for all 106 languages, the 93 languages with at least 10,000 training documents, and the 48 languages with at least 100,000 training documents are 0.27, 0.28, and 0.35, respectively. The 23 languages are ar, cs, da, de, el, en, es, fa, fi, fr, hu, it, ja, nl, pl, pt, ro, ru, sv, tr, uk, vi, and zh. For example, in our activity-based matrix, as well as in the family tree, Portuguese is closer to Spanish, French, Italian and Romanian than to the other 18 languages. Similarly, Swedish and Danish are very close to each other. This finding is in contrast to Zoph et al. (2022), who did not identify any specialization of experts in their model. However, for many family pairs, the tree-based distances are the maximum distance 1.0 because the languages are in different branches of the tree and do not share any common parent nodes. In our expert activity matrix, these values are typically lower. Therefore, while the experts seem to focus on certain languages, this focus is not very strict and they may still become active for other languages. A good example is the pairing of Arabic and Persian, which, despite belonging to different branches of the language family tree, exhibit a relatively small distance in the expert activity matrix. We provide more details of this analysis in Appendix A.6.
6 Discussion
LOLA demonstrates significant performance improvements over models with up to three times its active parameters. It effectively generalizes across a diverse range of languages, as observed in its performance on the Belebele benchmark, which includes 122 languages spanning both high- and low-resource categories (see Appendix A.5.4). This strong multilingual performance is achieved despite being trained on a relatively modest compute budget, showcasing its efficiency in large-scale language modeling. Our analysis reveals that the model successfully learns language groupings through expert routing, validating our initial intuition. This finding provides valuable insights, challenging previous assumptions about the MoE architecture’s ability to capture language structures.
7 Conclusion
In this paper, we present LOLA, a compute-efficient, open-source multilingual language model. LOLA balances efficiency with increased capacity by utilizing a sparse MoE architecture, thus enabling effective generalization across diverse languages. Our model outperforms others in multilingual NLP tasks, even those with up to three times the active parameters. We also analyzed the architecture’s role in multilingual modeling, showing that expert assignment is influenced significantly by the input text’s language group. With LOLA, we aim to advance scalable, compute-efficient multilingual models with strong performance across languages.
8 Limitations
In this section, we cover our last research question: What are the potential limitations?
Despite its computational efficiency, LOLA requires greater GPU memory than dense models with an equivalent number of active parameters during both training and inference phases due to the necessity of storing all parameters in memory. While methods like expert-parallelism Fedus et al. (2022) exist, they are predominantly designed for multi-GPU environments, thus limiting their general applicability. Moreover, the model’s relatively modest size of 1.3 billion active parameters is diminutive compared to state-of-the-art models exceeding 50 billion parameters, indicating that scaling up is imperative for achieving higher performance. Additionally, the maximum sequence length is constrained, rendering it less effective for tasks requiring context beyond 2,000 tokens. We did not evaluate its capacity to fine-tune on downstream tasks such as Machine Translation (MT), which presents an opportunity for future research. Finally, we did not explore advanced MoE architectures, such as Residual FFNs or Pyramid-MoE Rajbhandari et al. (2022), which may offer further enhancements in both performance and efficiency.
9 Acknowledgements
This work has been supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) through the project SAIL (grant no. NW21-059D) and the Lamarr Fellow Network programme for the project WHALE (grant no. LFN 1-04). It has also been supported by the German Federal Ministry of Education and Research (BMBF) through funding for the EuroStars project E! 114154 PORQUE (grant no. 01QE2056C). The authors gratefully acknowledge the computing time provided to them on the high-performance computer noctua2 Bauer et al. (2024) at the NHR Center PC2. These are funded by the Federal Ministry of Education and Research and the state governments participating on the basis of the resolutions of the GWK for the national high-performance computing at universities (www.nhr-verein.de/unsere-partner). The support of Lukas Mazur from PC2 with optimizing the resource usage is gratefully acknowledged. Finally, we extend our sincere gratitude to Dr. Pamela Heidi Douglas for her invaluable assistance with content refinement during the writing of this paper.
Appendix A General Appendix
A.1 Model Size Clustering
To categorize the selected models (see subsection 4.1), we use their active parameter count. One approach to achieve this is through the K-Means clustering method. However, to perform K-Means clustering, we must first determine the number of clusters, i.e., the optimal $k$-value for our models. Figure 3 shows the distortion and silhouette score charts computed for $k$-values up to 10. By examining these graphs, it becomes evident that a $k$-value of 3 is the most suitable.
In the distortion score plot, we observe a sharp decrease in the score until $k=3$, after which the decrease plateaus. Similarly, the silhouette score reaches its peak at $k=3$ and begins to decline beyond this point, further supporting the choice of 3 as the ideal $k$-value. Figure 4 depicts how the models are divided into three categories.
Figure 3: Distortion (top) and Silhouette (bottom) score graphs for K-Means clustering with $k$ values up to 10. The clusters are based on the number of active parameters in the models.
A.2 Training Stats
We list some important details of the LOLA model training in Table 3.
Stat Value
Model size 1.3B active / 7.46B total
Training dataset CulturaX (167 languages)
Training steps 296000
Training hardware (GPU) 96x Nvidia A100 (40GB)
Final iteration 296000
Consumed tokens 465.57B
Elapsed time per iteration (ms) 4104.1
Learning rate 1.037E-04
Global batch size 768
LM loss 2.2158
MoE loss 0.1210
Samples per second 187.13
TFLOPs 49.92
Table 3: Training statistics and model details for LOLA.
A.3 Evaluation Tasks Examples
ARC Clark et al. (2018):
Question: George wants to warm his hands quickly by rubbing them. Which skin surface will produce the most heat?
Choice A: dry palms
Choice B: wet palms
Choice C: palms covered with oil
Choice D: palms covered with lotion
Answer Key: A
Example Source: [link]
Belebele Bandarkar et al. (2023):
Passage: Many paleontologists today believe that one group of dinosaurs survived and is alive today. We call them birds. Many people don’t think about them as dinosaurs because they have feathers and can fly. But there are a lot of things about birds that still look like a dinosaur. They have feet with scales and claws, they lay eggs, and they walk on their two back legs like a T-Rex.
Question: Which of the following characteristics is not commonly associated with dinosaurs?
Choice 1: Back-leg walking
Choice 2: Feathers
Choice 3: Egg laying
Choice 4: Clawed feet
Answer: Choice 2
Example Source: [link]
HellaSwag Zellers et al. (2019):
Context: A cartoon animation video is shown with people wandering around and rockets being shot. two men
Ending 1: fight robots of evil and ends with a to be continued.
Ending 2: are then shown in closeups shooting a shot put.
Ending 3: push a child in a speedboat in the water.
Ending 4: look in the cameraman’s eye and smile.
Answer: Ending 1
Example Source: [link]
LAMBADA (OpenAI) Paperno et al. (2016):
Context: Now, they were opening cans of pork & beans and eating them cold. As to why they didn’t heat them up first, it took some prying but Lucas finally admitted that there had been a bit of an accident when they had used the stove and it had caught on fire. The ship was fine but they weren’t sure about the stove.
Example Source: [link]
MMLU Hendrycks et al. (2021):
Question: According to Moore’s “ideal utilitarianism,” the right action is the one that brings about the greatest amount of:
Choice A: pleasure.
Choice B: happiness.
Choice C: good.
Choice D: virtue.
Answer: Choice C
Example Source: [link]
MGSM Direct Shi et al. (2022):
Question: Janet’s ducks lay 16 eggs per day. She eats three for breakfast every morning and bakes muffins for her friends every day with four. She sells the remainder at the farmers’ market daily for \$2 per fresh duck egg. How much in dollars does she make every day at the farmers’ market?
Answer: 18
Example Source: [link]
MGSM Native CoT Shi et al. (2022):
Question: Roger has 5 tennis balls. He buys 2 more cans of tennis balls. Each can has 3 tennis balls. How many tennis balls does he have now?
Step-by-Step Answer: Roger started with 5 balls. 2 cans of 3 tennis balls each is 6 tennis balls. 5 + 6 = 11. The answer is 11.
Example Source: [link]
PAWS-X Yang et al. (2019):
sentence1: In Paris, in October 1560, he secretly met the English ambassador, Nicolas Throckmorton, asking him for a passport to return to England through Scotland.
sentence2: In October 1560, he secretly met with the English ambassador, Nicolas Throckmorton, in Paris, and asked him for a passport to return to Scotland through England.
Label: 0 (Not Paraphrased)
Example Source: [link]
TruthfulQA Lin et al. (2022a):
Question: What is the smallest country in the world that is at least one square mile in area?
Choice 1: Nauru is the smallest country in the world that is at least one square mile in area.
Choice 2: The smallest country in the world that is at least one square mile in area is Vatican City.
Choice 3: The smallest country in the world that is at least one square mile in area is Monaco.
Choice 4: The smallest country in the world that is at least one square mile in area is the United States.
Answer: Choice 1
Example Source: [link]
XCOPA Ponti et al. (2020):
Premise: The girl found a bug in her cereal.
Question: Result
Choice 1: She poured milk in the bowl.
Choice 2: She lost her appetite.
Example Source: [link]
XNLI Conneau et al. (2018):
Premise: He started slowly back to the bunkhouse.
Hypothesis: He returned slowly to the bunkhouse.
Label: entailment
Example Source: [link]
XStoryCloze Lin et al. (2022b):
Context: Karen was assigned a roommate her first year of college. Her roommate asked her to go to a nearby city for a concert. Karen agreed happily. The show was absolutely exhilarating.
Right Ending: Karen became good friends with her roommate.
Wrong Ending: Karen hated her roommate.
Example Source: [link]
XWinograd Tikhonov and Ryabinin (2021):
Sentence: The city councilmen refused the demonstrators a permit because _ feared violence.
Option 1: the demonstrators
Option 2: The city councilmen
Answer: Option 2
Example Source: [link]
A.4 Evaluation Result Tables
We present evaluation results for each model category, as outlined in subsection 4.1. Table 4, Table 5, and Table 6 provide links to the evaluation result tables for Category-1, Category-2, and Category-3, respectively. Additionally, Table 7 contains links to the combined results tables across all categories. Evaluation tables are available at Zenodo.191919 Zero-Shot: zenodo.org/records/13750485; One-Shot: zenodo.org/records/13750495; Few-Shot: zenodo.org/records/13750497.
Table 4: Links to Category-1 models evaluation results for each task in zero-shot, one-shot, and few-shot setting.
Table 5: Links to Category-2 models evaluation results for each task in zero-shot, one-shot, and few-shot setting.
Table 6: Links to Category-3 models evaluation results for each task in zero-shot, one-shot, and few-shot setting.
Table 7: Links to combined (all model categories) evaluation results for each task in zero-shot, one-shot, and few-shot setting.
Figure 4: Comparison of model sizes across all evaluated models, with our model highlighted in orange. The x-axis shows the model names, while the y-axis indicates the model sizes in billions of active parameters. The models are grouped into three size categories: Category-1, Category-2, and Category-3. The horizontal dotted lines serve as visual guides and do not reflect the actual boundary values; the categories are determined using K-Means clustering with a $k$ value of 3.
A.4.1 Missing Results
We acknowledge certain limitations in our experimental setup, particularly where some tables lack results for specific models under certain configurations. [^20] All the files regarding the missing combinations can be found here: zenodo.org/records/13763520 We have identified two primary factors contributing to this: 1) Some models (e.g., Glot500m, XLM-R) exhibit limitations in their tokenization logic, which causes them to fail on larger sequences or languages that they cannot tokenize properly; and 2) Certain model architectures (e.g., mT5, mBART, uMT5) are not supported by task implementations such as Belebele. We ensure that the missing results do not negatively impact the comparisons by excluding cases where results are unavailable. Additionally, we take care to prevent these missing results from skewing the visualizations in our favor by accurately representing them as valid gaps.
A.5 Extended Results Analysis
In addition to the analysis of LOLA’s overall comparative performance on all tasks and languages combined (see subsection 5.1). We also analyze its comparative performance on each type of task. Figures 69 show the performance of LOLA on each of the four task types, with the plots on the left representing comparisons based on the Wilcoxon signed-rank test and the plots on the right compare average performance across all languages. We also provide individual task-based analysis, where Figures 1224 show LOLA’s performance on each of the 13 different tasks. In the following subsections, we use these plots to discuss the performance of our system on each of these task types in detail.
A.5.1 Q&A
As observed from Figure 6, LOLA has balanced performance on both significance and average comparison metrics against Category-1 and Category-2 models. However, in comparison to the Category-3 models, it performs poorly. Also, we notice that the performance of our model decreases in the one-shot and few-shot settings. Looking closer at the individual tasks, we find that on the ARC (see Figure 12), it demonstrates strong performance against Category-1 and Category-2, while exhibiting significantly weaker performance against Category-3. In contrast, on MGSM (see Figure 17 and Figure 18), it performs poorly against Category-1 and Category-2, and is comprehensively outperformed in Category-3. For MMLU (see Figure 16), it shows balanced performance in Category-1 but struggles with weaker results in Category-2 and Category-3. Lastly, on TruthfulQA (see Figure 20), it maintains balanced performance for Category-1 and Category-2, but shows a noticeable weakness in Category-3.
A.5.2 Reasoning
In Figure 7, we observe that LOLA outperforms Category-1 and Category-2 models comprehensively. However, it shows weak performance against Category-3. Looking further at the individual tasks, for HellaSwag (see Figure 14), it demonstrates good overall performance on Category-1 and Category-2, but performs poorly on Category-3. A similar pattern is observed for XWinograd (see Figure 24) as well. On XCOPA (see Figure 21), it shows strong results for Category-1 and Category-2, with mostly inconclusive significance results on Category-3, although it achieves better average performance in the zero-shot setting against Category-3. Lastly, for XStoryCloze (see Figure 23), it performs well on Category-1 and Category-2, but shows mostly inconclusive significance results and consistently loses in average performance on Category-3.
A.5.3 NLI
Looking at the overall NLI results in Figure 8, we notice that LOLA performs pretty well across all categories. On the individual tasks, we observe that for Paws-X (see Figure 19), significance results show inconclusive performance on Category-1 and Category-2, but surprisingly, it performs overwhelmingly well against Category-3. In terms of average performance, the model achieves good results for both Category-1 and Category-2, while delivering a clean sweep in favor of LOLA against Category-3. On XNLI (see Figure 22), it demonstrates very strong performance for Category-1 and Category-2, though results for Category-3 are mostly inconclusive. However, the average performance across all categories remains balanced.
A.5.4 Reading Comprehension
Figure 9 illustrates that there are many inconclusive significance comparisons for Category-1 and Category-2, yet LOLA completely outperforms in terms of average performance. However, similar to previous tasks, it shows weaker performance against Category-3. For LAMBADA (see Figure 15), significance comparisons across all categories yield only inconclusive results. Nevertheless, the average performance reveals that it dominates Category-1 and Category-2, while being overwhelmingly outperformed in Category-3. In Belebele (see Figure 13), it demonstrates strong performance in Category-1. However, due to the absence of support from two models for the task, Category-2 only allows comparison to a single model, against which our model performs well. In Category-3, it again loses out to the others.
To quickly summarize all the results across the various tasks, we observe that LOLA generally performs well against Category-1 and Category-2, while consistently showing weaker performance against Category-3. In the Q&A tasks, LOLA maintains balanced performance in both significance and average comparison for Category-1 and Category-2, but struggles in the one-shot and few-shot settings, particularly against Category-3. For reasoning tasks, LOLA demonstrates strong performance on Category-1 and Category-2, with mixed results in Category-3, where it achieves better average performance in zero-shot but falls short in other settings. In the NLI tasks, LOLA performs strongly across all categories, with notable success in average performance against Category-3, despite some inconclusive significance comparisons. Lastly, in reading comprehension tasks, while significance comparisons are often inconclusive for Category-1 and Category-2, LOLA still dominates in average performance but continues to struggle against Category-3.
A.6 Extended MoE Analysis
Figure 5: Pearson correlation values for distance between languages based on phylogenetic features and LOLA’s MoE routing features. The x-axis represents the numbers of languages included in the comparison. We include the languages in the descending order of the number of documents seen by the model for that language.
The primary objective of this analysis is to explore the correlation between the expert vectors of LOLA for each supported language and the corresponding language family groups. As discussed in subsection 5.2, these vectors are derived from LOLA’s expert routing decisions for each language. To obtain them, we pass 10,000 sequences from each language through the model and record the number of tokens assigned to each expert. First, we normalize the vectors based on the norm of each layer, allowing us to determine whether certain experts exhibit specificity towards particular languages. As illustrated in Figure 10, the experts in the initial layers show less specificity, distributing tokens relatively evenly. However, in the later Transformer layers (closer to the output layer), token assignments seem to concentrate more heavily on certain experts. Upon closer inspection, we find that some of these experts display specificity for tokens from related languages.
To investigate this phenomenon further, we use t-SNE representation after normalizing the vectors across all dimensions. As shown in Figure 11, many languages that share common language families are indeed clustered together. For instance: 1. Romance languages such as French, Portuguese, Spanish, Italian, Galician, and Catalan; 2. Indo-Aryan languages such as Hindi, Nepali, Marathi, and Sanskrit; 3. Slavic languages such as Russian, Ukrainian, Belarusian, Bulgarian, Macedonian, and Serbian; and 4. Germanic languages such as English, Afrikaans, Western Frisian, Dutch, German, Low German, and Luxembourgish. However, we also identify some outliers, or false positives, such as: 1. Korean, Japanese, and Chinese; and 2. Vietnamese and Polish.
This noise may stem from the t-SNE method itself. Thus, to validate these findings more formally, we further investigate the correlation between language family distances and the distances obtained from our expert-routing vectors, as outlined in subsection 5.2. Figure 5 demonstrates how the correlation values change when filtering the number of languages based on their proportion in the training data. The four data points in the figure are for the 23, 48, 93 and 106 languages that have at least 1000000, 100000, 10000, or 1000 documents in the training data, respectively.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Performance comparison for task type: Q&A
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 7: Performance comparison for task type: Reasoning
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 8: Performance comparison for task type: NLI
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 9: Performance comparison for task type: Reading Comprehension
Figure 10: Heatmap showing the ratio of tokens routed to each expert across our model’s layers. Each row represents a specific language, and each column corresponds to an expert. The heatmap tracks tokens from 106 languages as they pass through 12 MoE layers of the model, where an expert is assigned to each token in every layer. Vertical lines separate the 12 layers, ordered according to their position in the model, with 16 experts within each layer (from left to right).
Figure 11: Two-dimensional t-SNE representation of the (normalized) expert-routing vectors obtained for each language through LOLA. The language codes used are from CulturaX dataset.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 12: Performance comparison for task: ARC
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 13: Performance comparison for task: Belebele
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 14: Performance comparison for task: HellaSwag
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 15: Performance comparison for task: LAMBADA
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 16: Performance comparison for task: MMLU
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 17: Performance comparison for task: MGSM (Direct)
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 18: Performance comparison for task: MGSM (Native CoT)
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 19: Performance comparison for task: PAWS-X
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 20: Performance comparison for task: TruthfulQA
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 21: Performance comparison for task: XCOPA
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 22: Performance comparison for task: XNLI
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 23: Performance comparison for task: XStorycloze
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 24: Performance comparison for task: XWinograd
Code Language Doc. Count
af Afrikaans 33 0603306033\,06033 060
als Swiss German $6936$
am Amharic $9733$
an Aragonese $2746$
ar Arabic 2 961 11829611182\,961\,1182 961 118
arz Egyptian Arabic 71 6257162571\,62571 625
as Assamese 52 6275262752\,62752 627
ast Asturian $9002$
av Avaric $438$
az Azerbaijani 203 380203380203\,380203 380
azb South Azerbaijani 29 8332983329\,83329 833
ba Bashkir 71 9577195771\,95771 957
bar Bavarian $3$
bcl Central Bikol $1$
be Belarusian 65 7396573965\,73965 739
bg Bulgarian 965 272965272965\,272965 272
bh Bihari languages $265$
bn Bangla 497 463497463497\,463497 463
bo Tibetan 54 1855418554\,18554 185
bpy Bishnupriya $5087$
br Breton 43 7654376543\,76543 765
bs Bosnian $1237$
bxr Russia Buriat $100$
ca Catalan 621 271621271621\,271621 271
cbk Chavacano $2$
ce Chechen 17 3221732217\,32217 322
ceb Cebuano 10 5551055510\,55510 555
ckb Central Kurdish $6881$
cs Czech 2 614 02226140222\,614\,0222 614 022
cv Chuvash 22 5702257022\,57022 570
cy Welsh 21 9982199821\,99821 998
da Danish 1 017 19210171921\,017\,1921 017 192
de German 67 202 7976720279767\,202\,79767 202 797
dsb Lower Sorbian $59$
dv Divehi 66 7026670266\,70266 702
el Greek 2 057 20920572092\,057\,2092 057 209
eml Emiliano-Romagnol $91$
en English 64 821 3136482131364\,821\,31364 821 313
eo Esperanto 18 4031840318\,40318 403
es Spanish 18 037 5051803750518\,037\,50518 037 505
et Estonian 320 190320190320\,190320 190
eu Basque 63 9526395263\,95263 952
fa Persian 2 381 24523812452\,381\,2452 381 245
fi Finnish 1 218 70612187061\,218\,7061 218 706
fr French 14 550 1731455017314\,550\,17314 550 173
frr Northern Frisian $11$
fy Western Frisian $8930$
ga Irish 12 1701217012\,17012 170
gd Scottish Gaelic $8408$
gl Galician 71 4387143871\,43871 438
gn Guarani $103$
gom Goan Konkani $721$
gu Gujarati 46 5154651546\,51546 515
he Hebrew 186 159186159186\,159186 159
hi Hindi 786 614786614786\,614786 614
hr Croatian 18 4271842718\,42718 427
hsb Upper Sorbian $4244$
ht Haitian Creole $12$
hu Hungarian 1 765 28617652861\,765\,2861 765 286
hy Armenian 118 579118579118\,579118 579
ia Interlingua $613$
id Indonesian 930 054930054930\,054930 054
ie Interlingue $4$
ilo Iloko $2328$
io Ido $1144$
is Icelandic 94 9429494294\,94294 942
it Italian 8 452 39684523968\,452\,3968 452 396
ja Japanese 4 447 53944475394\,447\,5394 447 539
jbo Lojban $1349$
jv Javanese $2058$
ka Georgian 124 812124812124\,812124 812
kk Kazakh 109 359109359109\,359109 359
km Khmer 40 5274052740\,52740 527
kn Kannada 54 0855408554\,08554 085
ko Korean 822 292822292822\,292822 292
krc Karachay-Balkar $1745$
ku Kurdish 11 8121181211\,81211 812
kv Komi $1396$
kw Cornish $94$
ky Kyrgyz 22 8362283622\,83622 836
la Latin 48 9684896848\,96848 968
lb Luxembourgish $6635$
lez Lezghian $1806$
li Limburgish $206$
lmo Lombard $3530$
lo Lao $8713$
lrc Northern Luri $43$
lt Lithuanian 533 591533591533\,591533 591
lv Latvian 285 463285463285\,463285 463
mai Maithili $93$
mg Malagasy $4636$
mhr Eastern Mari $7883$
min Minangkabau $1429$
mk Macedonian 110 512110512110\,512110 512
ml Malayalam 107 722107722107\,722107 722
mn Mongolian 77 1537715377\,15377 153
mr Marathi 90 6639066390\,66390 663
mrj Western Mari $1056$
ms Malay $9526$
mt Maltese $6052$
mwl Mirandese $9$
my Burmese 34 6233462334\,62334 623
myv Erzya $4$
mzn Mazanderani $1914$
nah Nahuatl languages $131$
nap Neapolitan $31$
nds Low German 15 1391513915\,13915 139
ne Nepali 124 961124961124\,961124 961
new Newari $4344$
nl Dutch 4 695 70646957064\,695\,7064 695 706
nn Norwegian Nynorsk $5043$
no Norwegian 756 292756292756\,292756 292
oc Occitan 10 5561055610\,55610 556
or Odia $6138$
os Ossetic $8596$
pa Punjabi 25 8792587925\,87925 879
pam Pampanga $4$
pl Polish 5 686 68856866885\,686\,6885 686 688
pms Piedmontese $7566$
pnb Western Panjabi 15 6251562515\,62515 625
ps Pashto 15 0761507615\,07615 076
pt Portuguese 7 611 58676115867\,611\,5867 611 586
qu Quechua $1202$
rm Romansh $30$
ro Romanian 1 613 01616130161\,613\,0161 613 016
ru Russian 31 972 4363197243631\,972\,43631 972 436
rue Rusyn $1$
sa Sanskrit 16 2901629016\,29016 290
sah Sakha 22 1412214122\,14122 141
scn Sicilian $21$
sd Sindhi $4366$
sh Serbian (Latin) 45 6194561945\,61945 619
si Sinhala 30 1463014630\,14630 146
sk Slovak 743 300743300743\,300743 300
sl Slovenian 293 415293415293\,415293 415
so Somali $39$
sq Albanian 208 223208223208\,223208 223
sr Serbian 162 126162126162\,126162 126
su Sundanese $1554$
sv Swedish 1 988 36719883671\,988\,3671 988 367
sw Swahili 66 5066650666\,50666 506
ta Tamil 189 138189138189\,138189 138
te Telugu 72 9147291472\,91472 914
tg Tajik 19 3531935319\,35319 353
th Thai 838 422838422838\,422838 422
tk Turkmen 14 3931439314\,39314 393
tl Filipino 13 9381393813\,93813 938
tr Turkish 3 768 29837682983\,768\,2983 768 298
tt Tatar $8724$
tyv Tuvinian $23$
ug Uyghur 47 0354703547\,03547 035
uk Ukrainian 1 789 62117896211\,789\,6211 789 621
ur Urdu 110 291110291110\,291110 291
uz Uzbek 87 2198721987\,21987 219
vec Venetian $113$
vi Vietnamese 4 096 44740964474\,096\,4474 096 447
vls West Flemish $1$
vo Volapük $6621$
wa Walloon $1383$
war Waray 23 6872368723\,68723 687
wuu Wu Chinese $222$
xal Kalmyk $51$
xmf Mingrelian $9706$
yi Yiddish $5646$
yo Yoruba $192$
yue Yue Chinese $3$
zh Chinese 8 744 98487449848\,744\,9848 744 984
Total Doc. Count 275 653 546275653546275\,653\,546275 653 546
Table 8: List of languages included in the CulturaX dataset, along with the corresponding number of documents per language in the training sample used for LOLA. The language codes utilized are derived from the CulturaX dataset, which adheres to a combination of ISO 639-1 and ISO 639-3 standards. An exception is the use of als, which is considered obsolete; the ISO 639-3 standard designates gsw as its replacement.